# OptiMUS-0.3: Using large language models to model and solve optimization problems at scale

### v2

@@ 4.4 Structure Detector Agent
OptiMUS also maintains a problem structure pool that enumerates combinatorial optimization problems with special structure that can be exploited in fast custom solvers (e.g., SAT, network flow, routing problems). For instance, the traveling salesman problem (TSP) can be solved much more efficiently using a specialized solver such as Concorde as opposed to general purpose MILP solvers (Cook et al. 2011). OptiMUS iterates through this pool of special problem structures given an initial natural language description of the problem. When the optimization problem falls into one of the problem types in the structure pool, OptiMUS identifies the particular problem structure and suggests using a customized solver via a notification in the OptiMUS webapp.
@@ 4.5 Advanced Optimization Coding Agent
For extremely large-scale optimization problems, optimization solvers are often embedded into a high-level optimization framework and called as a subroutine: examples include decomposition algorithms such as column generation, Bender’s decomposition, and cutting plane methods. To leverage variable and constraint structure in this large-scale context, the subproblem solver must invoke an advanced solver interface (for example, call-backs, model attribute query and analysis). While we leave the problem of developing a fully automated system that can perform advanced decomposition algorithms to future work, we have added several features to OptiMUS that advance this vision. OptiMUS implements a dedicated coding agent that exploits advanced solver functionality such as callbacks, and can iteratively call a solver within a simple variable or constraint sifting scheme (also known as delayed constraint generation). Neither of these modules require explicitly reformulating the problem (e.g., generating a pricing problem within column generation), but allow for better computational performance than a naive implementation.
Similar to the structure detector agent, the optimization coding agent maintains a series of template prompts for this functionality. Within each prompt, the LLM is provided with the purpose of the template. We illustrate this functionality with an application to sifting, a simple version of constraint or column generation, in Section 4.5.1. The advanced coding agent runs during the Code Clauses stage of OptiMUS (see Fig. 1).
@@ 4.5.1 Sifting
Figure 6: Left: constraint sifting prompt. Right: performance plot of solving to 5% gap.
@@ 5 Experiments
In this section, we conduct a comprehensive evaluation of OptiMUS-0.3. We showcase the superior performance of OptiMUS-0.3 on the NLP4LP dataset [^3] In the tests, we assume that the parameter names are given and identified in the problem description., highlighting its strengths and weaknesses. An ablation study demonstrates the impact of different system components on our results, and a sensitivity analysis probes the internal dynamics of OptiMUS-0.3. We conclude this section by identifying failure cases and potential areas for further improvement.
@@ 5.1 Overall Performance
To evaluate the overall performance of OptiMUS, we compare it with standard prompting, Reflexion, and Chain-of-Experts (CoE) (Shinn et al. 2023, Xiao et al. 2023). Reflexion is the highest-performing general-purpose framework and CoE is the state-of-the-art method for natural-language optimization modeling. We also include results from the intial conference publication of this work, which we denote as OptiMUS-0.2. Unless otherwise stated, we use GPT-4o (OpenAI 2023) as the LLM for OptiMUS and all baseline prompting methods.
@@
Three main metrics have been used to assess the accuracy of LP modeling tools in the literature: accuracy, compilation error (CE) rate, and runtime error (RE) rate. However, a method can generate irrelevant short code that runs, or fix runtime and compilation errors by completely removing relevant sections of the code. Hence, we only compare the models’ accuracy. Accuracy is defined as the number of instances correctly solved. An instance is considered as correctly solved only if 1) the code runs successfully; 2) the optimal value is correct; 3) the optimal solution is correct. Optimal values and solutions are obtained from the dataset or by solving the problems manually. Since the output of the code does not necessarily match the format of the optimal solution in the dataset, we use an LLM to determine whether the solution generated by OptiMUS matches the optimal solution in the dataset.
@@
Results are presented in Table 3. For both easy and hard problems, OptiMUS-0.3 outperforms all other methods in all datasets by a large margin, beating the standard GPT-4o baseline by over 40% for hard problems. This remarkable performance improvement highlights the importance of modularity and structure compared to a single prompt to solve complex problems using LLMs. OptiMUS relies on structured code outputs (OpenAI 2024, LangChain 2024), which are currently incompatible with inference-time reasoning models, such as OpenAI’s o1-preview. As of this writing, this limitation prevented us from testing these models in our experiments.
@@ 5.2 Ablation Study
Table 3 shows the impact of the choice of LLM on the performance of OptiMUS. We evaluate the OptiMUS system with both a leading closed-source LLM, GPT-4o, as well as a weaker open-source alternative, LLaMa-3.1-70B-Instruct (Meta AI 2024). As expected, the performance of the OptiMUS-0.3 framework degrades when less capable LLMs, such as LLaMa-3.1-70B-Instruct, are used instead of GPT-4o, especially on hard modeling tasks. Optimization tasks need complex reasoning, and smaller models may be better suited for simpler tasks. However, the OptiMUS-0.3 framework with LLaMa-3.1-70B-Instruct outperforms GPT-4o on with previous state-of-the-art prompting strategies, including Reflexion, on easy problems. This result highlights that the modular design of OptiMUS-0.3 not only allows it to outperform other prompting strategies using the same model, but can also makes the system more robust to the quality of the underlying LLM, allowing weaker cheaper models to match the performance of stronger models on simple optimization modeling tasks.
Table 3 also shows the impact of different components of the OptiMUS framework on its performance. Debugging (i.e., running error correction on the output of the Coding stage) has the largest impact on the performance of OptiMUS, increasing its accuracy on hard problems by almost 50%. The debugging step, is particularly important for harder problems, given LLMs often make a few small mistakes when they generate code. Error correction also leads to modest improvements in the performance of OptiMUS. Running error correction after extracting clauses is particularly important for difficult problems. The LLM feedback step marginally improves the performance of the system. Note that we do not include the impact of the structure detection agent and advanced optimization coding agent as they do not impact the accuracy of the framework, but just the efficiency (i.e., runtime). For examples showcasing the impact of these components, see Appendix 11.
Fig. 7 (left) shows the relation between the performance and the number of debugging steps for different GPT-4o and LLaMa-3.1-70B-Instruct. In both models, the first few debugging steps improve performance, but further steps do not result in improvement.
@@ 5.3 Failure Cases
We analyze the most common reasons why OptiMUS fails, categorizing failure cases into the following groups (Fig. 7):
@@
Table 2: OptiMUS-0.3 outperforms the other methods with GPT-4o
Easy Hard
Standard 47.3% 33.2%
Reflexion 53.0% 42.6%
CoE 64.2% 49.2%
OptiMUS-0.2 78.8% 68%
OptiMUS-0.3 86.6% 73.7%
@@
GPT4-o 86.6% 73.7%
1111222233334444555566660020202020404040406060606080808080100100100100Number of Debugging IterationsPerformance (%)GPT4-o EasyGPT4-o HardLLaMa3-70B EasyLLaMa3-70B Hard
ExtractionModelingCoding00101010102020202030303030404040405050505060606060707070708080808090909090100100100100Failure TypeError rate (%)EasyHard

### v3

@@ 4.4 Structure Detection Agent
OptiMUS also maintains a problem structure pool that enumerates combinatorial optimization problems with special structure that can be exploited in fast custom solvers (e.g., SAT, network flow, routing problems). For instance, the traveling salesman problem (TSP) can be solved much more efficiently using a specialized solver such as Concorde as opposed to general-purpose MILP solvers (Cook et al. 2011). OptiMUS iterates through this pool of special problem structures given an initial natural language description of the problem. When the optimization problem falls into one of the problem types in the structure pool, OptiMUS identifies the particular problem structure and suggests using a customized solver via a notification in the OptiMUS webapp.
@@ 4.5 Advanced Optimization Coding Agent
For extremely large-scale optimization problems, optimization solvers are often embedded into a high-level optimization framework and called as a subroutine: examples include decomposition algorithms such as column generation, Benders decomposition, and cutting plane methods. To leverage variable and constraint structure in this large-scale context, the subproblem solver must invoke an advanced solver interface (for example, call-backs, model attribute query and analysis). While we leave the problem of developing a fully automated system that can perform advanced decomposition algorithms to future work, we have added several features to OptiMUS that advance this vision. OptiMUS implements a dedicated coding agent that exploits advanced solver functionality such as callbacks, and can iteratively call a solver within a simple variable or constraint sifting scheme (also known as delayed constraint generation). Neither of these modules requires explicitly reformulating the problem (e.g., generating a pricing problem within column generation), but allow for better computational performance than a naive implementation.
Similar to the structure detection agent, the optimization coding agent maintains a series of template prompts for this functionality. Within each prompt, the LLM is provided with the purpose of the template. We illustrate this functionality with an application to sifting, a simple version of constraint or column generation, in Section 4.5.1. The advanced coding agent runs during the Code Clauses stage of OptiMUS (see Fig. 1).
@@ 4.5.1 Sifting
Figure 6: Left: constraint sifting prompt. Right: performance plot for solving to a 5% gap.
@@ 5 Experiments
In this section, we conduct a comprehensive evaluation of OptiMUS-0.3. We showcase the superior performance of OptiMUS-0.3 on both the NLP4LP dataset and two existing benchmark datasets for optimization modeling highlighting its strengths and weaknesses. An ablation study demonstrates the impact of different system components on our results, and a sensitivity analysis probes the internal dynamics of OptiMUS-0.3. We conclude this section by identifying failure cases and potential areas for further improvement.
@@ 5.1 Overall Performance
To evaluate the overall performance of OptiMUS, we compare it with standard prompting, Reflexion, and Chain-of-Experts (CoE) (Shinn et al. 2023, Xiao et al. 2023). Reflexion is the highest-performing general-purpose framework, and CoE is a recently proposed agentic framework for natural-language optimization modeling. We include self-reported results for two state-of-the-art auto-formulation methods based on fine-tuning LLMs specifically for optimization modeling: ORLM (Tang et al. 2024), and LLMOPT (Jiang et al. 2024). We also include results from the initial conference publication of this work, which we denote as OptiMUS-0.2. We report results for the OptiMUS-0.3 framework with both GPT-4o (OpenAI 2023) and o1 (Jaech et al. 2024), an LLM that automatically includes more advanced reasoning at the expense of longer inference times. Due to the cost of running o1, we report performance with standard prompting from Lu et al. (2025) and only run the OptiMUS framework on datasets where o1 alone does not already achieve near-perfect accuracy. In our experiments, we use the default API parameters for all LLM calls.
Evaluation Metrics and Datasets
@@
Three main metrics have been used to assess the accuracy of LP modeling tools in the literature: accuracy, compilation error (CE) rate, and runtime error (RE) rate. However, a method can generate an irrelevant short code that runs, or fix runtime and compilation errors by completely removing relevant sections of the code. Hence, we only compare the models’ accuracy. Accuracy is defined as the number of instances correctly solved. An instance is considered as correctly solved only if 1) the code runs successfully, 2) the optimal value is correct, and 3) the optimal solution is correct. Optimal values and solutions are obtained from the dataset or by solving the problems manually. Since the output of the code does not necessarily match the format of the optimal solution in the dataset, we use an LLM to determine whether the solution generated by OptiMUS matches the optimal solution in the dataset. In addition to the NLP4LP dataset introduced in this paper, we include results for two existing datasets for optimization modeling: NL4OPT (Ramamonjison et al 2023) and IndustryOR (Tang et al. 2024). The latter dataset includes 100 real-world case studies from eight different industries. In the tests, we assume that the parameter names are given and identified in the problem description.
Table 2: Accuracy of auto-formulation methods on benchmark datasets. indicates performance taken directly from original paper, and blanks indicate no results reported on the dataset.
LLM NL4OPT NLP4LP IndustryOR
Methods based on direct prompting
Standard GPT-4o 47.3% 33.2% 28.0%
Standard o1 >> 95% 68.8% 44.0%
Reflexion GPT-4o 53.0% 42.6%
Methods based on fine-tuning LLMs
LLMOPT Qwen1.5-14B 93.0% 83.8% 46.0%
ORLM Deepseek-Math 86.5% 72.9% 38.0%
Methods based on agentic frameworks
CoE GPT-4o 64.2% 49.2%
OptiMUS-0.2 GPT-4o 78.8% 68.0%
OptiMUS-0.3 GPT-4o 86.6% 73.7% 37.0%
OptiMUS-0.3 o1 80.6% 46.0%
@@
Results are presented in Table 2. OptiMUS-0.3 outperforms all other direct prompting and agentic methods in all datasets by a large margin, beating the standard GPT-4o baseline by over 40% on the NLP4LP dataset. This improvement persists even with more advanced reasoning models such as o1. This impressive performance improvement highlights the importance of modularity and structure compared to a single prompt to solve complex problems using LLMs.
Remarkably, the OptiMUS framework also remains competitive with finetuning approaches, beating the performance of ORLM and remaining within 4% of LLMOPT. Note that fine-tuning based frameworks — which must be trained and deployed by the user, and updated via expensive retraining to take advantage of a new base model —represent a much more logistically complex approach compared to agentic frameworks that can be run with simple API calls to LLM providers and instantly improved by access to a new base model. We see that modular design of an agentic system can close the performance gap between direct prompting and finetuning, giving practitioners a simpler alternative to state-of-the-art systems. The improved performance of the OptiMUS-0.3 framework with o1 over GPT-4o also underscores a key benefit of an agentic framework: as more advanced models are released, the framework can be easily adjusted to run with the new models by substituting a simple API call whereas finetuning methods require repeating the costly finetuning procedure.
On Computation Time
In Figure 7 we investigate the running time of the OptiMUS framework across 85 randomly selected instances of NLP4LP using GPT-4o and along different stages of the pipeline. The OptiMUS-0.3 framework takes no more than 350 seconds (median time 108 seconds) to model and solve any problem instance. In contrast, recent work highlighted that human subject matter experts take on average approximately 150 minutes for similar tasks (Tang et al. 2024). Coding and debugging represent the most time-consuming parts of the OptiMUS pipeline. Smaller LLM agents specialized to coding (e.g., codex) could further improve the computation time of the framework.
Figure 7: Distribution of running time (s) for different stages in the OptiMUS framework.
On Stochasticity in LLM Outputs
LLM outputs are well-known to be probabilistic, meaning that calling an LLM multiple times can lead to different outputs. To evaluate the impact of this stochasticity on the performance of OptiMUS-0.3 we re-ran 10 hard instances of the NLP4LP dataset over five different random seeds. Across all instances, the result (i.e., success or failure) remained consistent across all 5 random seeds, providing strong evidence that OptiMUS is robust to this source of stochasticity. We hypothesize that the EC modules help mitigate performance variability arising from stochastic LLM outputs. Conversely, we do not observe improved performance for multiple samples of the LLM output. One interpretation is that for hard instances, new strategies, and not simply repeated samples, are needed to improve performance.
On Calibration
A key feature of the OptiMUS-0.3 system is the confidence-based LLM feedback detailed in Section 4.3.2. To evaluate the calibration of the LLM-based confidence scores we sampled 40 confidence scores from the clause modeling module and enlisted a doctoral student in operations research to manually annotate whether each clause was correct. Of the 40 clauses, the LLM responded with a confidence score of 5/5 in 28/40 of the instances and was correct 100100% of the time, and a score of less than 5 in 12/40 instances and was correct 91.791.7% of the time. This provides promising evidence that the LLM-based confidence scores can effectively guide a user to identify errors. Exploring more advanced mechanisms for eliciting calibrated confidence scores (Xiong et al. 2023) is a promising direction for future work.
@@ 5.2 Ablation Study
Table 3 shows the impact of the choice of LLM on the performance of OptiMUS, broken down by its performance on hard and easy instances of the NLP4LP dataset. We evaluate the OptiMUS system with a leading closed-source LLM, GPT-4o, a reasoning model, o1, as well as a weaker open-source alternative, LLaMa-3.1-70B-Instruct (Meta AI 2024). As expected, the performance of the OptiMUS-0.3 framework degrades when less capable LLMs, such as LLaMa-3.1-70B-Instruct, are used instead of GPT-4o or o1, especially on hard modeling tasks. Optimization tasks need complex reasoning, and smaller models may be better suited for simpler tasks. However, the OptiMUS-0.3 framework with LLaMa-3.1-70B-Instruct outperforms GPT-4o with previous state-of-the-art prompting strategies, including Reflexion, on easy problems. We see that OptiMUS-0.3 is more robust to the quality of the underlying LLM, allowing weaker cheaper models to match the performance of stronger models on simple optimization modeling tasks.
Table 3 also shows the impact of different components of the OptiMUS framework on its performance. Debugging (i.e., running error correction on the output of the Coding stage) has the largest impact on the performance of OptiMUS, increasing its accuracy on hard problems by almost 50%. The debugging step, is particularly important for harder problems, given LLMs often make a few small mistakes when they generate code. Error correction also leads to modest improvements in the performance of OptiMUS. Running error correction after extracting clauses is particularly important for difficult problems. Table 4 further breaks down the impact of error correction methods on correct and incorrect items (i.e., constraints). In both constraint extraction and modeling, EC modules are able to fix a large fraction of errors without modifying most correct items. The LLM feedback step also marginally improves the performance of the system. Note that we do not include the impact of the structure detection agent and advanced optimization coding agent as they do not impact the accuracy of the framework, but just the efficiency (i.e., runtime). For examples showcasing the impact of these components, see Appendix 10.
Fig. 8 (left) shows the relation between the performance and the number of debugging steps for OptiMUS with both GPT-4o and LLaMa-3.1-70B-Instruct. In both models, the first few debugging steps improve performance, but further steps do not result in improvement.
@@ 5.3 Failure Cases
We next analyze the most common reasons why OptiMUS fails. We categorized failure cases manually via a qualitative coding process based in grounded theory (Charmaz 2006). We followed an open-coding process to analyze instances on which OptiMUS failed and tag the type of error. These initial codes were then grouped together into the following coarser categories:
@@
GPT-4o 86.6% 73.7%
o1 80.6%
11223344556602020404060608080100100Number of Debugging IterationsPerformance (%)GPT4-o EasyGPT4-o HardLLaMa3-70B EasyLLaMa3-70B Hard
ExtractionModelingCoding010102020Failure TypeError rate (%)EasyHard
@@ 7 Conclusion
Acknowledgments
AA, WG, CL, and MU gratefully acknowledge support from the National Science Foundation (NSF) Award IIS-2233762, the Office of Naval Research (ONR) Awards N000142212825, N000142412306, and N000142312203, the Alfred P. Sloan Foundation, and from IBM Research as a founding member of Stanford Institute for Human-centered Artificial Intelligence (HAI). HB thanks the Aker Scholarship Foundation for financial support. ST gratefully acknowledges support from the Air Force Office of Scientific Research (AFOSR) under Grant FA9550-23-1-0251 and in part by the Office of Naval Research under Grant N00014-24-1-2164. This manuscript was improved by feedback from anonymous referees.